OpenStudy (anonymous):

Find the x- and y-intercepts for the equation. Please show all of your work 2x – 4y = 1

OpenStudy (anonymous):

x- and y-intercepts are respectivly where y=0 and x=0 on the graoh. So put this values to find them

OpenStudy (anonymous):

So 2(0)-4(0)=1 is my solution?

OpenStudy (anonymous):

but do it 1 at a time. y=0 for x-intcpt 2x-4*0 =1 and later the other

OpenStudy (anonymous):

Ok, so 2x -4*0 =1 And 2*0-4x =1 are my equations correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you son much. How do i give medals for your help?

OpenStudy (anonymous):

push blue button "Best Response"

OpenStudy (anonymous):

Thank you again for your help. finishing up my homework now.

OpenStudy (anonymous):

you wellcome